OpenStudy (anonymous):

can someone explain and show there work on how to do this problem (3t^2)^3(2t^0)^-3

OpenStudy (anonymous):

\[ (3t^2)^3 = 3^3(t^2)^3=27t^{2\times 3}=27t^6 \]

OpenStudy (anonymous):

That is how you would do part of it. Try to do t he whole thing on your own.

OpenStudy (anonymous):

would the problem result in a fraction @wio

OpenStudy (anonymous):

Yes, likely it will.

OpenStudy (anonymous):

27t^6/8

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

okay thanks